OpenStudy (anonymous):

A Ferris wheel car moves from point C to point D on the circle shown below: Circle A is shown with a central angle marked 42 degrees and the diameter marked 30 feet. What is the arc length the car traveled, to the nearest hundredth?

OpenStudy (anonymous):

7.91 feet 8.32 feet 10.99 feet 11.89 feet @welshfella

OpenStudy (calculusxy):

Arc Length Formula: \[\frac{ central \angle }{ 360 degrees} \times circumference \]

OpenStudy (calculusxy):

To find the circumference: \[C = 2\pi r\]

OpenStudy (anonymous):

Ok so c=2pi(15)=94.24

OpenStudy (calculusxy):

Yes then use that with the \[\frac{ central \angle }{ 360 degrees }\]

OpenStudy (anonymous):

so \[\frac{ 42 }{ 360 }\] x 94.24?

OpenStudy (calculusxy):

Yes

OpenStudy (anonymous):

so that would be 10.99 which is the third answer choice! thanks so much!!

OpenStudy (anonymous):

Do you mind if i ask a couple more questions?

OpenStudy (calculusxy):

I'll try to help

OpenStudy (anonymous):

ok i'll set it up as a new question so i can medal u again